OpenStudy (anonymous):

Hi can any one tell me Why gradient vector is orthogonal to tangent ? with an example

OpenStudy (amistre64):

F(x,y) = x - y(x) Fx = 1 Fy = -1y' grad(F) = <1,-1y'> notice that F(x,y) = x-y(x) is equalvalent to: y = x, with a tangent of y'=1 at any x

OpenStudy (amistre64):

the gradient vector is a normal vector, and a normal vector is always perp to a tangent vector

OpenStudy (anonymous):

. that's a sphere and it is the easiest way to see how they perp together